A damped harmonic oscillator loses 8 percent of its mechanical energy per cycle. (a) By what percentage does its frequency differ from the natural frequency f0 = (1/2?)?k/m?
Answer #1

For damped harmonic motion A= Ao e st Here r= 6/2m wiek z I wo 11- 11 J m After one period energy decreases to 8%. E = 0.92Eoin cq. o becomes wl-wo a 1 1 - 1 1 1 I 1-187 12 - : wo wo w-wo, a 2.205x10-5 % difference (wh- wo ) lov = (2-205x105 100 of

Hope this helps....